### v1

# Fair Play in the Newsroom: Actor-Based Filtering Gender Discrimination in Text Corpora

Abstract
Large language models are increasingly shaping digital communication, yet their outputs often reflect structural gender imbalances that originate from their training data. This paper presents an extended actor-level pipeline for detecting and mitigating gender discrimination in large-scale text corpora. Building on prior work in discourse-aware fairness analysis, we introduce new actor-level metrics that capture asymmetries in sentiment, syntactic agency, and quotation styles. The pipeline supports both diagnostic corpus analysis and exclusion-based balancing, enabling the construction of fairer corpora. We apply our approach to the taz2024full corpus of German newspaper articles from 1980 to 2024, demonstrating substantial improvements in gender balance across multiple linguistic dimensions. Our results show that while surface-level asymmetries can be mitigated through filtering and rebalancing, subtler forms of bias persist, particularly in sentiment and framing. We release the tools and reports to support further research in discourse-based fairness auditing and equitable corpus construction.
Fair Play in the Newsroom:
Actor-Based Filtering Gender Discrimination in Text Corpora
Stefanie Urchs1,2, Veronika Thurner1, Matthias Aßenmacher2,3, Christian Heumann2, Stephanie Thiemichen1 1Faculty for Computer Science and Mathematics, Hochschule München University of Applied Sciences, 2Department of Statistics, LMU Munich, 3Munich Center for Machine Learning (MCML), LMU Munich Correspondence: stefanie.urchs@hm.edu
1 Introduction
Large Language Models (LLM) are increasingly integrated into everyday digital services, from search engines and translation tools to conversational agents and content recommendation systems. Despite their impressive capabilities, LLMs are known to perpetuate and even amplify harmful societal biases, including gender stereotypes, occupational hierarchies, and asymmetric visibility of social groups Torrielli (2025); Armstrong et al. (2024); Siddique et al. (2024); Jeoung et al. (2023). These patterns are not inherent to the models’ architecture itself (Vaswani et al., 2017) but are learned from the data on which they are trained. As training data plays a central role in shaping the representational landscape of LLMs’ outputs, critical analysis of large-scale text corpora becomes a key step towards mitigating downstream harms. Without an understanding of the social structures embedded in textual data, efforts to ensure fairness and accountability in LLMs risk addressing symptoms rather than causes.
Previous work by Urchs et al. (2025) introduced a linguistically grounded pipeline to detect gender discrimination in German newspaper texts through actor-level discourse analysis. Their approach identifies named actors and analyses how they are represented in terms of visibility and associated language, drawing on concepts such as nomination and predication from critical discourse studies.
Building on this foundation, we present an extended pipeline that enables both fine-grained fairness auditing and corpus-level discrimination reduction. Our contribution is fourfold:
  1. 1.
    We introduce novel actor-level discrimination markers, including syntactic roles, quote attribution, and sentiment bias.
  2. 2.
    We enhance interpretability through structured, human-readable reports that enable qualitative and diachronic analysis.
  3. 3.
    We propose a method for generating gender-balanced corpus recommendations by excluding disproportionately discriminatory texts.
  4. 4.
    We release the entire pipeline as open-source software to support transparency, reproducibility, and collaborative development.
This paper offers tools and insights for creating fairer NLP systems by revealing how social groups are represented in text. We combine discourse-informed fairness analysis with scalable corpus processing to enable actor-level discrimination detection and targeted corpus balancing.
2 Related Work and Conceptual Background
Detecting gender discrimination in text requires an interdisciplinary foundation that integrates perspectives from linguistics, gender studies, and computer science. In this section, we first outline our conceptual understanding of gender and linguistic discrimination. We then review existing approaches to computational discrimination detection. Finally, we introduce the actor-level discrimination detection pipeline by Urchs et al. (2025), which forms the foundation of our work.
2.1 Gender and Linguistic Discrimination
In this work, we adopt a differentiated understanding of gender and discrimination that draws from linguistic discourse analysis, gender studies, and computational fairness research.
Gender is treated as a socially constructed and co-constructed identity, rather than a fixed biological or grammatical category. While linguistic gender relates to grammatical rules and biological sex is often seen as binary and immutable, we work with the concept of social gender, which can be fluid, non-binary, and shaped through interaction and recognition. This perspective foregrounds the performative and contextual nature of gender, acknowledging its entanglement with social norms, symbolic power, and habitualised practices West and Zimmerman (1987); Ainsworth (2015); Konishi (1993); Kramer (2020); Devinney et al. (2022). However, due to methodological constraints and the structure of the German language, the empirical analysis remains limited to the binary spectrum.
Discrimination, in contrast to bias or fairness, is understood here as a social effect: It is the observable outcome of differential treatment based on protected attributes such as gender. Drawing on Reisigl (2017) and the functional model by Graumann and Wintermantel (2007), we conceptualise linguistic discrimination as an act or process that disadvantages (or favours) individuals through patterns in language. This includes acts of naming (nomination) and describing (predication) actors in ways that reflect or reinforce social hierarchies. Such acts need not be intentional, but they may also arise from the habitual reproduction of dominant norms in public discourse. This framework stands in contrast to many machine learning–based approaches, where bias often refers to statistical imbalances in data or model performance, and fairness is operationalised via metrics such as demographic parity or equal opportunity Blodgett et al. (2020); Caton and Haas (2024). While these frameworks are useful for quantifying group disparities, they tend to reduce complex social identities to binary categories, thereby risking the oversight of contextual, structural, and discursive forms of inequality.
2.2 Computational Discrimination Detection
Discrimination in text is not limited to overtly offensive statements but often emerges from subtle and structural patterns of language use. Computational discrimination analysis aims to make such patterns visible by identifying systematic disparities in how individuals or groups are represented in text. In computer science, discrimination is often formalised via the concept of fairness and operationalised using statistical metrics such as demographic parity, equalised odds, or individual fairness Mehrabi et al. (2021). These metrics are well-suited for classification tasks and support scalability and reproducibility. However, they also reduce social categories to fixed, binary attributes and abstract away from contextual and structural forms of inequality Blodgett et al. (2020).
When applied to text, this paradigm produces approaches that focus primarily on hate speech detection, sentiment disparity, or stereotyping, often based on keyword lists or supervised classification. Although these methods yield valuable insights, they rarely address how discrimination is embedded in discourse or how it is distributed across different individuals within a text. Instead, most existing systems operate at a text-level granularity, assigning a global label such as “discriminatory” or “non-discriminatory” to entire documents.
In contrast, Urchs et al. (2025) focus on discrimination at the actor level. By identifying individual actors and examining how they are named (nomination) and described (predication), their approach reveals asymmetries in representation within the same text. This enables a more granular and discourse-aware analysis of discrimination in language.
2.3 Actor-Level Discrimination Detection Pipeline
Our pipeline builds upon prior work by Urchs et al. (2024, 2025). The first paper introduces actor-based fairness analysis in isolated texts using a modular pipeline that combines information extraction with linguistic discourse analysis. Specifically, it detects gender discrimination on the actor level by identifying the nomination and the predication of individual actors. The pipeline extracts actors via named entity recognition (NER), resolves pronouns through coreference resolution, and stores all actor references, including names, titles, and generic forms, in a structured knowledge base. For each actor, all sentences containing them are extracted and analysed for sentiment, gender-coded language, and linguistic framing. Discrimination metrics are computed per actor and summarised into a discrimination report. The report includes a range of linguistic and structural metrics designed to capture different facets of gendered representation:
  • Actor counts: Number of distinct male-, female-, and undefined-coded actors per text.
  • Mention counts: Total number of pronoun or name-based references per gender group.
  • Sentiment: Average sentiment score of all predications linked to each actor or gender group.
  • Gender-coded language: Count of feminine-coded and masculine-coded terms in predications, based on lexicons from Gaucher et al. (2011).
Their second paper (Urchs et al., 2025) scales this analysis to a large newspaper corpus (taz2024full) with over 1.8 million articles published between 1980 and 2024. It adapts the actor-level pipeline for German, replacing the English sentiment model with a BERT-based classifier trained on German news, and introduces additional markers for gender-neutral language and generic masculine usage. The analysis aggregates actor-based metrics by year, enabling the study of historical shifts in gender representation and framing.
Beyond the metrics introduced in the earlier paper, the taz2024full version adds:
  • Generic masculine detection: Flags texts using the German generic masculine form.
  • Gender-neutral language detection: Identifies inclusive writing styles such as gender colons or stars (e.g., Lehrer:innen).
  • PMI adjectives: Extracts the ten adjectives with the highest Pointwise Mutual Information (PMI) per actor, providing insights into recurring descriptive patterns.
  • Yearly aggregation: Metrics are aggregated per year to enable longitudinal analysis of shifts in gendered representation and framing.
  • Yearly report generation: All extracted metrics are compiled into a structured, human-readable report for each year.
However, the approach remains purely descriptive and diagnostic: No mechanism is implemented for correcting or filtering the corpus based on the findings. Our work extends the work from both papers substantially in depth and scope. We expand the actor-level analysis with new discrimination detection metrics and integrate the pipeline into a two-stage exclusion framework.
3 The Extended Actor-Centred Pipeline
We extend the original actor-centred pipeline, improving both the granularity of the analysis and its ability to support fairer corpus construction.
Building on insights from systemic functional linguistics Halliday (2004) and critical discourse analysis Reisigl (2017), the pipeline now incorporates syntactic role annotation to capture how subject and object positions contribute to gendered representations. In discourse, actors in subject roles typically perform actions, while those in object positions are acted upon. Tracking these roles across pronoun groups helps reveal patterns of agency and passivity that are central to linguistic discrimination. To deepen the understanding of the gendered representation, we track whether actors are referred to by their name or just by their pronouns. We refined the quote attribution by distinguishing direct from indirect speech using punctuation and reporting verbs. Furthermore, zooming in on how the actors’ utterances are presented gives insights into how active they are in the text. Direct quotes indicate a higher degree of activity than indirect ones. We enrich the framing context by extending pointwise mutual information (PMI) calculations to include not only adjectives but also verbs and nouns. This expansion highlights thematic associations and role-specific language tied to gendered actors. Human-readable reports are redesigned to improve structure, interpretability, and accessibility. All results are presented in a structured plain-text report, with separate sections for summary statistics, syntactic roles, sentiment, and lexical framing. Breakdowns on a pronoun-group basis and PMI tables are aligned and consistently formatted, making the output interpretable even for non-technical readers.
Finally, the pipeline supports a two-step corpus filtering mechanism: Articles exhibiting strong internal gender asymmetries are flagged using multiple indicators, and a subsequent balancing step adjusts overall gender ratios in the dataset. This process yields a curated corpus suitable for training language models with reduced gender discrimination.
The full pipeline code is available on our git repository https://github.com/Ognatai/corpus_balancing.
4 Pipeline Application: Discrimination Analysis and Corpus Balancing
This section outlines how we implemented the extended actor-level pipeline in two consecutive stages.
4.1 Stage 1: Discrimination Analysis Across the Full Corpus
For the initial discrimination analysis, we build on the pipeline proposed by Urchs et al. (2025) (cf. §2.3). We extend this approach by incorporating a set of additional linguistic and structural metrics that allow for a more fine-grained assessment of gendered representation and framing. These metrics are extracted at the actor level and aggregated per document and year:
  • Mentions: Captures how often an actor is referred to by name (e.g., Angela Merkel) or pronoun (e.g., she, he), enabling analysis of individuation, visibility, and referential strategies.
  • Syntactic roles: Counts how often the actor appears as grammatical subject (nsubj) or object (obj), providing a proxy for discursive agency and passivity.
  • Quotation style: Differentiates between direct and indirect speech attributions, reflecting variation in narrative presence and framing control.
  • Top 10 PMI terms: Lists the nouns and verbs most strongly associated with the actor based on Pointwise Mutual Information (PMI), offering insight into typical roles, actions, and semantic contexts.
The full taz2024full corpus is analysed both at the yearly level and aggregated in total. We also introduce an improved yearly reporting format that organises both existing and newly introduced metrics in a more accessible structure. An example of this yearly report (for 2023) can be found in Appendix A.
4.2 Stage 2: Article Filtering and Corpus Balancing
To reach the endeavour of constructing a more balanced corpus and substantially reducing the impact of gender-discriminating articles, we introduce a multi-stage filtering pipeline. In contrast to stage 1, which analyses the corpus on a yearly basis, this stage operates on the entire dataset to enable global exclusion and balancing decisions.
Step 1: Full Corpus Analysis
The full corpus is processed using the pipeline from Stage 1 (cf. §2.3). Instead of aggregating metrics per year, actor-level knowledge bases and article-level metadata are saved for each article. These intermediate files are reused in the subsequent steps to minimise redundant executions of pipeline steps. After processing, a histogram is generated to visualise the distribution of gender ratios across articles (cf. Figure 5). Each article is assigned two values: the percentage of mentions and the percentage of actors associated with she/her pronouns. The left subplot displays the mention-weighted distribution, and the right subplot shows the actor-weighted distribution. Both histograms use percentage values ranging from 0 (only he/him references) to 100 (only she/her references), with each bar representing the proportion of articles falling into that range.
Step 2: Text-Level Exclusion
An interactive filtering interface allows the user to define how many of the four heuristic exclusion criteria (see list below) must be met to classify a text as discriminatory. The default threshold is two out of four. The criteria are:
  • Sentiment disparity: A large gap in average sentiment scores between female- and male-coded actors.
  • Grammatical role asymmetry: A strong difference in subject-to-object ratios between pronoun groups.
  • Quote attribution imbalance: A pronounced disparity in direct versus indirect quotes.
  • Naming versus pronoun imbalance: Disproportionate usage of named versus pronominal references between genders.
Each criterion is calculated using Laplace-smoothed ratios to avoid instability possibly originating from low-frequency counts. The user may customise the threshold for each criterion at runtime. By default, a text exhibits significant imbalance if the sentiment gap exceeds 0.3, or if the difference in subject/object roles, quote attribution, or naming/pronoun usage exceeds 0.5. These thresholds correspond to cases where one gender group is at least twice as prominent as the other in a specific framing dimension. A lower threshold is used for sentiment because sentiment values tend to cluster around neutral and vary within a narrower range; even small differences may indicate meaningful affective bias. Overall, the chosen defaults strike a balance between interpretability and selectivity, capturing strongly biased texts while preserving as much of the corpus as possible. Articles that trigger the user-defined number of flags are excluded from the corpus. Exclusion decisions are logged in a structured exclusion file and visualised through an updated gender ratio distribution plot (cf. Figure 6).
Step 3: Corpus-Level Balancing
After text-level filtering, we apply corpus-level balancing based on the relative contribution of each article to the overall gender ratio. The user sets an equilibrium range for actor and mention ratios between female- and male-coded references. The default range is $[0.75,1.25]$, allowing for up to 25% deviation in either direction. Articles are iteratively excluded based on their contribution to the global imbalance until both ratios fall within the specified interval. A visualisation of the resulting gender ratio distribution is generated to illustrate the effects of balancing (cf. Figure 7.) All excluded articles are recorded in a structured exclusion file for transparency and reproducibility.
Step 4: Corpus Reconstruction
All excluded article IDs from both filtering stages are consolidated and used to construct a new balanced corpus. Articles are removed directly from the original JSON files, and the revised dataset is saved to disk.
5 Corpus-Balancing of taz2024full
We use the taz2024full corpus Urchs et al. (2025), comprising over 1.8 million articles from the German left-leaning newspaper taz (1980–2024), previously used to analyse gender representation. In the unfiltered taz2024full corpus, we detect female- and male-coded actors in 1,834,018 articles. Gender representation is clearly imbalanced: Men dominate across both mention frequency and actor counts (Figure 1). These differences are not only quantitative but also reflected in patterns of agency and authority.
Figure 1: Percentage of male- and female-coded references over time before filtering. Fluctuations in the early years are due to the small number of articles available from the 1980s, which can lead to disproportionate weight for single-gender articles.
Figure 2 shows that male actors are more frequently quoted directly, while female actors appear more often in indirect quotes. Direct quotations often attribute more authority to the speaker and allow them to appear in active, public-facing roles. Indirect quotes, by contrast, reduce visibility and typically appear in paraphrased or backgrounded contexts.
Figure 2: Distribution of quotation styles by gender before filtering. Early-year fluctuations are again attributable to low article counts.
Figure 3 further underlines this pattern. Men occur more frequently in subject positions, while women appear comparatively more often as objects. Subject roles in grammar typically denote agency and narrative control, whereas object positions signal reduced agency and passivity within the sentence structure.
Figure 3: Distribution of syntactic roles by gender before filtering. Early-year fluctuations are again attributable to low article counts.
Sentiment analysis reveals a consistent gap: The average sentiment towards female-coded actors is more negative in almost all years (Figure 4). While sentiment values are generally close to neutral, the persistent divergence reinforces the broader imbalance.
Figure 4: Average sentiment associated with male- and female-coded actors before filtering.
The early-year spikes in Figures 1, 2, and 3 are artefacts of low data density. In years with very few articles or actor mentions, a single gender may dominate, resulting in sharp fluctuations that do not reflect structural bias but instead data sparsity. We retain these years for completeness, but advise caution when interpreting early data points and refrain from doing this ourselves.
Finally, the overall distribution of gender representation per article (Figure 5) is highly polarised: many articles reference either only male-coded or only female-coded actors. This reinforces the need for corpus-level balancing, as it shows that article-level imbalance is not just a matter of aggregate statistics but of individual article composition.
Figure 5: Distribution of gender ratios across articles before filtering.
During the first text-level filtering step, we exclude 279,772 articles based on four framing asymmetries: sentiment gap, quote imbalance, subject/object ratio, and representation. We use the default values described in Section 4.2. The updated diagnostic view shows reduced polarisation, and the gender ratio distribution shifts closer to balance (Figure 6).
Figure 6: Distribution of gender ratios across articles after text-level filtering.
Subsequently, we apply corpus-level balancing, excluding an additional 17,815 articles to bring the overall actor and mention ratios into the (default) target range $[0.75,1.25]$. Compared to the filtered corpus, the distribution is more centred and less polarised, indicating that articles with extreme gender dominance were successfully downsampled to achieve a more balanced overall representation. (Figure 7).
Figure 7: Distribution of gender ratios across articles after corpus-level balancing.
In the final corpus, gender representation is nearly balanced across both mentions and actor counts (Figure 8). This balance is not just a numerical artefact of article exclusion, but reflects a more even distribution across time and actor types. Compared to the original corpus (cf. Figure 1), the trajectories of female- and male-coded representation converge, with both actor and mention lines approaching parity. Importantly, the gender crossing point around 2018 persists, indicating that key corpus dynamics remain intact after balancing.
Figure 8: Percentage of male- and female-coded references over time after all filtering and balancing steps.
Figure 9 shows that the quotation gap is reduced compared to Figure 2. Women now appear in direct speech more frequently than before, while men remain slightly more dominant in indirect speech. The remaining asymmetry may reflect residual stylistic preferences in journalistic practice or shifts in topic domains rather than systematic bias. Overall, this pattern reflects a notable improvement in discursive agency, with women quoted more often in their own words.
Figure 9: Proportion of direct and indirect quotations by gender after full exclusion.
Sentiment remains slightly more negative for women (cf. Figure 10), but the gap remains below the exclusion threshold of 0.3. This persistence may signal that subtle evaluative framing is harder to eliminate via structural balancing. It also suggests that sentiment operates on a different level than grammatical or referential bias, possibly rooted more deeply in lexical associations or broader discourse conventions.
Figure 10: Average sentiment towards she/her and he/him actors over time after full exclusion.
Figure 11 reveals a modest shift in syntactic agency. The proportion of women in subject positions increases slightly, while object roles are more evenly distributed. Although men still dominate in subject roles, the difference is reduced compared to the original corpus. Specifically, the subject-role gap between men and women decreases from approximately 30 to 5 percentage points after filtering. This suggests that our multi-step exclusion process leads to subtle improvements in how grammatical agency is distributed between genders.
Figure 11: Distribution of syntactic roles after full exclusion.
Taken together, these results demonstrate the effectiveness of our multi-stage filtering and balancing strategy in mitigating structural gender asymmetries in the corpus. By comparing before and after exclusion, we observe clear improvements across key dimensions of representation and framing.
Referential parity is achieved: male- and female-coded actors now appear in comparable proportions across both mentions and actor counts. Discursive patterns also show notable shifts. Women are quoted directly more often, and syntactic agency is redistributed more evenly, with a marked reduction in the subject-object gap. These improvements suggest that our approach meaningfully alters how gendered actors are positioned within the narrative structure of the corpus.
However, some subtle forms of bias remain. The sentiment gap persists below the exclusion threshold, indicating that evaluative language is harder to correct through structural rebalancing alone. Such residual asymmetries may reflect deeper, more diffuse biases embedded in lexical or thematic choices rather than in sentence-level grammar.
Overall, the final corpus provides a significantly more equitable foundation for downstream NLP applications and for critical media analysis. By combining large-scale actor-level auditing with targeted corpus interventions, our approach offers a concrete path towards fairer data curation in practice.
6 Conclusion and Future Work
In this paper, we presented an extended actor-level pipeline for detecting and mitigating gender discrimination in large-scale text corpora. Building on prior work, we introduced new metrics that capture asymmetries in syntactic roles, quote attribution, and sentiment framing. We enhanced the interpretability of the results through structured reports and implemented a two-stage filtering mechanism that enables the construction of gender-balanced corpora.
Our application of the pipeline to the taz2024full corpus demonstrates that gender imbalances in representation and framing are both measurable and correctable to a significant extent. The resulting corpus shows improved balance across multiple linguistic dimensions and serves as a more equitable foundation for downstream tasks such as language model training or critical media studies.
Nevertheless, some forms of discrimination, particularly those tied to sentiment and more implicit discourse structures, persist despite structural balancing. This indicates that not all bias can be addressed through surface-level interventions alone. Future work should therefore explore complementary strategies such as employing context-aware language models for deeper semantic analysis, developing targeted debiasing methods to address persistent framing asymmetries, and integrating intersectional attributes such as race, age, or class. In addition, extending actor categories beyond the gender binary would enable the inclusion of non-binary and gender-diverse identities, allowing for a more comprehensive understanding of representational fairness.
More broadly, we advocate for the integration of discourse-aware methods into standard corpus construction workflows. Understanding how social groups are framed in language is a necessary prerequisite for designing fairer NLP systems, and our pipeline offers a scalable, modular, and linguistically grounded way to do so.
Use of AI
The authors are not native English speakers; therefore, ChatGPT and Grammarly were used to assist with writing English in this work. ChatGPT was also used to assist with coding.
Limitations
While our approach enables corpus-level balancing based on measurable framing asymmetries, it is not without limitations. First, the exclusion-based strategy necessarily reduces corpus size and diversity, potentially eliminating valuable content alongside discriminatory texts. Second, the method operates on surface-level linguistic signals and cannot fully account for subtler or context-dependent forms of bias, such as irony, framing through omission, or topic selection. Third, the balancing relies on binary gender classification, which excludes non-binary identities and reinforces a gender dichotomy that our conceptual framework otherwise seeks to challenge. Fourth, the analysis is limited to texts that contain clearly identifiable actors and sufficient gender cues, primarily via pronoun resolution. Articles without identifiable pronouns or mentions of coreferent actors are excluded from the discrimination analysis altogether, leading to incomplete coverage. Finally, the impact-aware exclusion method is sensitive to threshold settings and metric selection, which may affect outcomes in ways that are not always transparent. These limitations highlight the need for complementary strategies, such as counterfactual augmentation, contextual bias detection, or narrative-level analysis, to address more complex and nuanced forms of representational inequality.
Ethical Considerations
Our work is grounded in the belief that fairness in NLP requires not only technical interventions but also critical reflection on the social impact of language technologies. By analysing how gendered actors are represented and framed in text, we aim to make structural inequalities visible and address them at the level of data design. At the same time, we recognise that fairness cannot be reduced to numerical balance. The act of filtering texts, however principled, entails normative decisions about which content is considered discriminatory and which is preserved. This introduces risks of over-correction, loss of valuable context, and the potential erasure of complex identities. Furthermore, our reliance on binary gender resolution excludes non-binary, intersex, and gender-nonconforming individuals, reinforcing the very simplifications we seek to critique. We consider this a significant ethical limitation and prioritise extending our methods to support more inclusive representations in future work. Finally, while we aim to mitigate bias in training data, we stress that ethical responsibility must also extend to model architectures, deployment contexts, and the broader socio-technical systems in which NLP tools are embedded.
Acknowledgments
Matthias Aßenmacher is funded with funds from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI - grant number 460037581. Responsibility for the contents of this publication lies with the authors.
Appendix
Appendix A Corpus Report 2023
Report for the year 2023
AGGREGATED TOTALS (all texts)
Total Texts: 10019
Texts with Actors: 10019
Uses Gender Neutral Language (Docs): 107
Generic Masculine Usage (Docs): 8081
Metric she/her he/him overall
Pronoun Distribution: 6892 9194 16086
Mentions by Pronoun: 35595 56044 91639
Named Mentions: 22544 36047 58591
Pronoun Mentions: 13051 19997 33048
Subject Roles: 18625 30303 48928
Object Roles: 1119 1540 2659
Direct Quotes: 6501 10588 17089
Indirect Quotes: 2529 4215 6744
Feminine-coded Words: 4251 6066 10317
Masculine-coded Words: 2870 4764 7634
Sentiment: -0.01 -0.01 -0.01
Named Mentions (% of all mentions): 38.5 61.5
Pronoun Mentions (% of all mentions): 39.5 60.5
Subject Roles (% of known roles): 38.1 61.9
Object Roles (% of known roles): 42.1 57.9
Direct Quotes (% of quotes): 38.0 62.0
Indirect Quotes (% of quotes): 37.5 62.5
STATISTICS (per text)
Metric Mean Median Std Dev
Pronouns (Resolved) (She/Her) 0.69 1.00 0.83
Mentions (By Pronoun) (She/Her) 3.55 2.00 7.22
Feminine Coded Words (By Pronoun) (She/Her) 0.42 0.00 1.18
Masculine Coded Words (By Pronoun) (She/Her) 0.29 0.00 0.77
Named Mentions (Sum Over Actors) (She/Her) 2.25 1.00 5.65
Pronoun Mentions (Sum Over Actors) (She/Her) 1.30 1.00 2.29
Subject Roles (She/Her) 1.86 0.00 3.78
Object Roles (She/Her) 0.11 0.00 0.43
Direct Quotes (She/Her) 0.65 0.00 1.46
Indirect Quotes (She/Her) 0.25 0.00 0.72
Pronouns (Resolved) (He/Him) 0.92 1.00 0.91
Mentions (By Pronoun) (He/Him) 5.59 3.00 9.77
Feminine Coded Words (By Pronoun) (He/Him) 0.61 0.00 1.36
Masculine Coded Words (By Pronoun) (He/Him) 0.48 0.00 1.05
Named Mentions (Sum Over Actors) (He/Him) 3.60 1.00 7.75
Pronoun Mentions (Sum Over Actors) (He/Him) 2.00 1.00 3.02
Subject Roles (He/Him) 3.02 2.00 5.01
Object Roles (He/Him) 0.15 0.00 0.53
Direct Quotes (He/Him) 1.06 0.00 1.96
Indirect Quotes (He/Him) 0.42 0.00 0.97
Mean Sentiment (All) -0.02 0.00 0.10
Total Actors 1.61 1.00 1.04
Total Mentions 9.15 5.00 12.10
Total Feminine Coded Words 1.03 0.00 1.83
Total Masculine Coded Words 0.76 0.00 1.28
Uses Gender-Neutral Language 0.01 0.00 0.10
Generic Masculine 0.81 1.00 0.40
TOP PMI ADJECTIVES
Most frequent adjectives associated with each pronoun group.
Rank ALL she/her he/him
1 letzten (414.00) letzten (154.00) letzten (269.00)
2 russischen (272.00) junge (130.00) russischen (195.00)
3 deutschen (260.00) berliner (101.00) deutschen (171.00)
4 berliner (231.00) deutschen (97.00) politische (142.00)
5 junge (212.00) deutsche (97.00) ukrainische (137.00)
6 nächsten (212.00) russischen (81.00) politischen (135.00)
7 politische (212.00) nächsten (80.00) berliner (134.00)
8 deutsche (208.00) politischen (80.00) nächsten (133.00)
9 politischen (205.00) politische (74.00) ukrainischen(117.00)
10 ukrainische (178.00) jungen (71.00) russische (113.00)
TOP PMI NOUNS
Most frequent nouns associated with each pronoun group.
Rank ALL she/her he/him
1 menschen (588.00) menschen (311.00) menschen (315.00)
2 frau (353.00) frau (234.00) präsident (289.00)
3 präsident (328.00) frauen (163.00) mann (210.00)
4 leben (312.00) leben (140.00) partei (185.00)
5 mann (280.00) mutter (128.00) leben (182.00)
6 partei (268.00) kinder (109.00) land (164.00)
7 land (238.00) tochter (107.00) frau (147.00)
8 frauen (210.00) geschichte (101.00) sohn (135.00)
9 stadt (209.00) mann (100.00) stadt (135.00)
10 regierung (208.00) anfang (100.00) mittwoch (126.00)
TOP PMI VERBS
Most frequent verbs associated with each pronoun group.
Rank ALL she/her he/him
1 erzählt (671.00) erzählt (331.00) erzählt (368.00)
2 steht (495.00) steht (199.00) steht (324.00)
3 sieht (449.00) erklärt (180.00) sieht (315.00)
4 erklärt (428.00) lassen (167.00) erklärt (269.00)
5 lassen (359.00) sieht (163.00) erklärte (243.00)
6 erklärte (346.00) sehen (147.00) spricht (228.00)
7 spricht (341.00) zeigt (139.00) lassen (205.00)
8 zeigt (302.00) spricht (139.00) sprach (199.00)
9 weiß (289.00) lebt (127.00) zeigt (190.00)
10 hält (286.00) sagen (125.00) weiß (188.00)